OpenStudy (anonymous):

A recent study of 200 nurses found that of 125 female nurses, 56 had bachelor's degrees; and of 75 male nurses, 34 had bachelor's degrees. If a nurse is selected at random, nd the probability that the nurse is: (a) A female nurse with a bachelor's degree. (b) A male nurse. (c) A male nurse with a bachelor's degree.

OpenStudy (anonymous):

are these correct? a. 7/25 b. 3/8 c. 17/100

OpenStudy (ddcamp):

a)56/200 = 7/25 b)75/200 = 3/8 c)34/200 = 17/100 Yup, looks like you got them all right!